# USP: Unified Self-Supervised Pretraining for Image Generation and Understanding

### v1

@@ Abstract
Recent studies have highlighted the interplay between diffusion models and representation learning. Intermediate representations from diffusion models can be leveraged for downstream visual tasks, while self-supervised vision models can enhance the convergence and generation quality of diffusion models. However, transferring pretrained weights from vision models to diffusion models is challenging due to input mismatches and the use of latent spaces. To address these challenges, we propose Unified Self-supervised Pretraining (USP), a framework that initializes diffusion models via masked latent modeling in a Variational Autoencoder (VAE) latent space. USP achieves comparable performance in understanding tasks while significantly improving the convergence speed and generation quality of diffusion models. Our code will be publicly available at https://github.com/cxxgtxy/USP.
@@ 1 Introduction
  • Unified Pretraining Framework. Inspired by the success of the pretraining-fine-tuning paradigm in image understanding, we propose USP, a unified pretraining framework that synergizes image comprehension and diffusion-based image generation. Upon completion of pretraining, learned representations are readily transferable to downstream tasks via weight initialization.
  • Decoupled Pretraining from Downstream Tasks. To decouple pretraining from the heterogeneous optimization objectives of downstream tasks, we propose masked feature modeling within the latent space of VAEs. This approach enables robust representation learning in an unsupervised manner, eliminating the need for labeled data and facilitating fast training.
  • Efficiency and Versatility. Our approach, through meticulous weight initialization, significantly enhances the performance of two transformer-based diffusion models: DiT and SiT. It achieves superior results compared to their baselines—trained for 7M steps—in just 600K and 150K steps, respectively, yielding speedups of 11.7$\times$ and 46.6$\times$. Importantly, it maintains strong representation capabilities for image recognition, underscoring its versatility. Additionally, it incurs no extra training cost or memory overhead, ensuring high efficiency and scalability.
@@ 2.2 Model Architecture
Given an input image $I\in H\times W\times 3$, its latent feature $I_{vae}$ is obtained through a frozen VAE. For example, using the SD-VAE [71] with $H=W=224$, we obtain a compressed feature $I_{vae}\in\frac{H}{8}\times\frac{W}{8}\times 4$. This latent feature is then divided into non-overlapping patches using 2$\times$2 convolutions (referred to as PatchConv), resulting in patches of size $\frac{H}{16}\times\frac{W}{16}\times h$, comparable to those used in [27]. A class token is added to the image tokens and enhanced with their sine-cosine position encodings. These patches are randomly masked with a specified mask ratio $m$. The unmasked patches are first processed by a standard Vision Transformer (ViT) encoder [27] with $h$ hidden dimensions, and then concatenated with the masked patches as input to an asymmetric vision decoder for reconstruction. After pretraining, the decoder is discarded, and only the ViT encoder and PatchConv are retained. Finally, the weights of the pretrained model are used to initialize downstream tasks such as classification, semantic segmentation, and generation. Therefore, this approach doesn’t introduce extra either training cost or memory overhead for downstream tasks.
Data augmentation and normalization. We adopt a weak data augmentation strategy, specifically RandomHorizontalFlip. This approach offers two key advantages: 1) it reduces the data augmentation gap between pretraining and d generation; 2) it allows for the use of caches to accelerate training. We utilize the normalization scheme of the VAE, instead of the widely used ImageNet normalization. Although this modification deviates from zero mean and unit variance, it ensures consistency in the data distribution between pretraining and fine-tuning.
@@ 2.3 Initialization Adaption
For image recognition tasks, the pretrained weights can be seamlessly inherited without additional effort. The class token is reused for classification. Specifically, we augment the model by incorporating a frozen VAE encoder and adjusting the patch projection size to align with the pretrained model architecture, ensuring a smooth transition and compatibility with downstream recognition tasks.
For generative tasks, we carefully adjust the network architecture to ensure consistency between the pretrained ViT model and the initialized DiT/SiT models. Specifically, we introduce slight modifications to the AdaLN-Zero layer normalization scheme. We reintroduce trainable bias $\beta$ and scale $\gamma$ parameters to align with the corresponding pretrained backbone. This adjustment represents a negligible increase in the total number of trainable parameters of DiT/SiT. Given that our model is pretrained at a resolution of 224$\times$224, we employ bicubic interpolation to upscale the positional embeddings from 224$\times$224 to 256$\times$256 for ImageNet generation tasks at the 256$\times$256 resolution. The class token is directly abandoned.
@@ 3.1 Pretraining
To ensure a fair comparison with MAE [39], we pretrain our models at a resolution of 224$\times$224, although experiments indicate that higher performance can be achieved with larger resolutions. We utilize a mask ratio of 0.75 and apply patch-wise normalization loss. The VAE is kept frozen throughout this paper. To maintain comparable training efficiency with MAE, we use the VAE to generate image latent features and construct caches in an off-the-shelf manner. Note that we adopt a different image normalization setting, with mean=[0.5, 0.5, 0.5] and std=[0.5, 0.5, 0.5], instead of the default ImageNet normalization.
@@ 3.2 Image Generation
We validate our method on two transformer diffusion models: DiTs [66] (based on DDPM [42]) and SiTs [62] (based on flow models [53, 62]. We adhere to the default training settings [66, 62]. For SiTs, we use linear interpolants and velocity predictions. Evaluation metrics include Fréchet Inception Distance (FID [41]), Inception Score (IS [72]), Precision (Pre.), and Recall (Rec.)[46]. Evaluations are conducted on ImageNet 256$\times$256 using 50K samples, without classifier-free guidance (CFG).
@@ 3.2.1 Comparison Under the DiT Framework
Model Params Steps FID ()($\downarrow$)( ↓ ) IS ()($\downarrow$)( ↓ )
@@ 3.2.2 Comparison Under the SiT Framework
3.2.3 Comparison with Acceleration Methods.
@@ 3.2.4 Training Cost
REPA FT 850K 9.6 1904 1833 $\dagger$ 71
Table 4: Training cost measured on DiT-XL/2 (in H20 GPU days). $\dagger$: estimated from [65] based on our test, where 1A100=2H20.
@@ 3.4 Ablation Study
We perform ablation study in this section and provide additional ablation studies in Section A.1 (appendix).
@@
(f) Different VAEs.
@@
Table 7: Ablation experiments. Default settings are marked in grey. Additional ablations are provided in Section A.1.
@@ 3.5 Discussion
Working Mechanism. As illustrated in Figure 4, we compare the linear probing performance of DiT-XL/2 pre- and post-generation training, as well as SiT-XL/2 with and without REPA training. Our findings underscore that effective initialization significantly enhances DiT’s recognition capability, with earlier layers capturing high-level semantic features and later layers optimized for generation, aligning with prior observations [97]. In contrast, REPA employs handcrafted alignment layers, which consistently exhibit superior linear performance (e.g., the 8th layer in Figure 4). While these layers demonstrate strong performance, they may not achieve an optimal balance between recognition and generation tasks. Our approach diverges from this methodology by leveraging a carefully curated initialization strategy that enables the network to autonomously identify the most effective layer for linear classification. This self-adaptive mechanism ensures an optimal balance between the dual objectives of recognition and generation. Notably, after 400k iterations of generation training using our proposed method, the 20th layer emerges as the optimal layer for linear classification performance. We attribute the enhanced performance of our method, in part, to this innovative mechanism.
@@ 5 Conclusion
Appendix A More Experiments
A.1 More Ablation Studies
@@ C.2 Fully Tuning on ImageNet
The hyperparameter setting for fine-tuning in ImageNet is shown in 12.
@@ C.3 Image Generation on ImageNet
The hyperparameter setting for generation in ImageNet is shown in 12.

### v2

@@ Abstract
Recent studies have highlighted the interplay between diffusion models and representation learning. Intermediate representations from diffusion models can be leveraged for downstream visual tasks, while self-supervised vision models can enhance the convergence and generation quality of diffusion models. However, transferring pretrained weights from vision models to diffusion models is challenging due to input mismatches and the use of latent spaces. To address these challenges, we propose Unified Self-supervised Pretraining (USP), a framework that initializes diffusion models via masked latent modeling in a Variational Autoencoder (VAE) latent space. USP achieves comparable performance in understanding tasks while significantly improving the convergence speed and generation quality of diffusion models. Our code is available on https://github.com/AMAP-ML/USP.
@@ 1 Introduction
  • Inspired by the success of the pretraining-fine-tuning paradigm in image understanding, we propose USP, a unified pretraining framework that synergizes image comprehension and diffusion-based image generation.
  • To decouple pretraining from the heterogeneous optimization objectives of downstream tasks, we propose masked feature modeling within the latent space of VAEs. This approach enables robust representation learning in an unsupervised manner, eliminating the need for labeled data and facilitating fast training.
  • Our approach, through meticulous weight initialization, significantly enhances the performance of two transformer-based diffusion models: DiT and SiT. It achieves superior results compared to their baselines—trained for 7M steps—in just 600K and 150K steps, respectively, yielding speedups of 11.7$\times$ and 46.6$\times$. Importantly, it maintains strong representation capabilities for image recognition, underscoring its versatility. Additionally, it incurs no extra training cost or memory overhead, ensuring high efficiency and scalability. Moreover, our method is somewhat orthogonal to other acceleration methods such as REPA [97] and VAVAE [95].
@@ 2.2 Model Architecture
Given an input image $I\in\mathbb{R}^{H\times W\times 3}$, its latent feature $I_{vae}$ is obtained through a frozen VAE. For example, using the SD-VAE [71] with $H=W=224$, we obtain a compressed feature $I_{vae}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 4}$. This latent feature is then divided into non-overlapping patches using 2$\times$2 convolutions (referred to as PatchConv), resulting in patches of size $\frac{H}{16}\times\frac{W}{16}\times h$, comparable to those used in [27]. A class token is added to the image tokens and enhanced with their sine-cosine position encodings. These patches are randomly masked with a specified mask ratio $m$. The unmasked patches are first processed by a standard Vision Transformer (ViT) encoder [27] with $h$ hidden dimensions, and then concatenated with the masked patches as input to an asymmetric vision decoder for reconstruction. After pretraining, the decoder is discarded, and only the ViT encoder and PatchConv are retained. Finally, the weights of the pretrained model are used to initialize downstream tasks such as classification, semantic segmentation, and generation. Therefore, this approach doesn’t introduce extra either training cost or memory overhead for downstream tasks.
Data augmentation and normalization. We adopt a weak data augmentation strategy, specifically RandomHorizontalFlip. This approach offers two key advantages: 1) it reduces the data augmentation gap between pretraining and generation; 2) it allows for the use of caches to accelerate training. We utilize the normalization scheme of the VAE, instead of the widely used ImageNet normalization. Although this modification deviates from zero mean and unit variance, it ensures consistency in the data distribution between pretraining and fine-tuning.
@@ 2.3 Initialization Adaption
For image recognition tasks, the pretrained weights can be seamlessly inherited without additional effort. The class token is reused for classification.
For generative tasks, we carefully adjust the network architecture to ensure consistency between the pretrained ViT model and the initialized DiT/SiT models. Specifically, we introduce slight modifications to the AdaLN-Zero layer normalization scheme. We reintroduce trainable bias $\beta$ and scale $\gamma$ parameters to inherit the weights of the pretrained backbone fully. This adjustment represents a negligible increase in the total number of trainable parameters of DiT/SiT. Given that our model is pretrained at a resolution of 224$\times$224, we employ bicubic interpolation to upscale the positional embeddings from 224$\times$224 to 256$\times$256 for ImageNet generation tasks at the 256$\times$256 resolution. The class token is directly abandoned.
@@ 3.1 Pretraining
To ensure a fair comparison with MAE [39], we pretrain our models at a resolution of 224$\times$224, although experiments indicate that higher performance can be achieved with larger resolutions. We utilize a mask ratio of 0.75 and apply patch-wise normalization loss as MAE [39]. The VAE is kept frozen throughout this paper. To maintain comparable training efficiency with MAE, we use the VAE to generate image latent features and construct caches in an off-the-shelf manner. Note that we adopt a different image normalization setting, with mean=[0.5, 0.5, 0.5] and std=[0.5, 0.5, 0.5], instead of the default ImageNet normalization.
@@ 3.2 Image Generation
We validate our method on two transformer diffusion models: DiTs [66] (based on DDPM [42]) and SiTs [62] (based on flow models [53]. We adhere to the default training settings [66, 62]. For SiTs, we use linear interpolants and velocity predictions. Evaluation metrics include Fréchet Inception Distance (FID [41]), Inception Score (IS [72]), Precision (Pre.), and Recall (Rec.)[46]. Evaluations are conducted on ImageNet 256$\times$256 using 50K samples, without classifier-free guidance (CFG).
@@ 3.2.1 Comparison Under the DiT Framework
Model Params Steps FID ()($\downarrow$)( ↓ ) IS ()($\uparrow$)( ↑ )
@@ 3.2.2 Comparison Under the SiT Framework
3.2.3 Comparison with Acceleration Methods
@@
3.2.4 Orthogonality to External-Model-Based Methods
Although USP achieves significantly faster convergence for diffusion models without relying on any external models, we further verify its orthogonality to recent DINO-based acceleration methods [97, 95]. As shown in Table 3, combining USP with either VAVAE [95] or REPA [97] leads to even faster convergence than using any single method alone.
Model Params Steps FID ()($\downarrow$)( ↓ ) IS ()($\uparrow$)( ↑ )
SiT-XL/2 130M 400K 16.97 77.50
USP 130M 400K 7.38 127.96
REPA 130M 400K 7.9 122.6
USP + REPA 130M 400K 6.26 139.84
VAVAE 130M 64 Epochs 5.18/2.15 132.4/245.1
USP + VAVAE 130M 64 Epochs 4.2/1.81 144/261.0
Table 3: Results Combined with External-Model-Based Methods. : w CFG=10.0.
@@ 3.2.5 Training Cost
REPA FT 850K 9.6 1904 1833 71
Table 5: Training cost measured on DiT-XL/2 (in H20 GPU days). : estimated from [65] based on our test.
@@ 3.4 Ablation Study
We perform ablation study in this section and provide additional ablation studies in Section A (appendix).
@@
(f) Different VAEs. C: channels, S: down-sampling factor.
@@
Table 8: Ablation experiments. Default settings are marked in grey. Additional ablations are provided in Section A.
@@ 3.5 Discussion
Working Mechanism. As illustrated in Figure 4, we compare the linear probing performance of DiT-XL/2 pre- and post-generation training, as well as SiT-XL/2 with and without REPA training. Our findings underscore that effective initialization significantly enhances DiT’s recognition capability, with earlier layers capturing high-level semantic features and later layers optimized for generation, aligning with prior observations [97]. In contrast, REPA employs handcrafted alignment layers, which consistently exhibit superior linear performance (e.g., the 8th layer in Figure 4). While these layers demonstrate strong performance, they may not be optimal. Our approach diverges from this methodology by leveraging a carefully curated initialization strategy that enables the network to autonomously identify the most effective layer for linear classification. Notably, after 400k iterations of generation training using our proposed method, the 20th layer emerges as the optimal layer for linear classification performance. We attribute the enhanced performance of our method, in part, to this innovative mechanism.
@@ 5 Conclusion
Supplementary Material
Appendix A More Ablation Studies
@@
High-Resolution Results Table 10 demonstrates the effectiveness of USP at a higher resolution of 512×512. We initialize the downstream image generation task using weights obtained from pretraining with USP at 256 resolution and directly transfer them to the 512 resolution setting, adjusting the positional encodings via bilinear interpolation. The results confirm that USP maintains strong performance and transferability under higher-resolution settings.
Model Params Steps FID ()($\downarrow$)( ↓ ) IS ()($\uparrow$)( ↑ )
SiT-B/2 130M 400K 42.80 37.37
USP 130M 400K 33.89 45.03
Table 10: Results at 512×512 Resolution.
@@ C.2 Fully Tuning on ImageNet
The hyperparameter setting for fine-tuning in ImageNet is shown in 14.
@@ C.3 Image Generation on ImageNet
The hyperparameter setting for generation in ImageNet is shown in 14.